Arasoo is 4 years older than Peter.Ali is 2 years younger than Peter.If the sum of their ages is 47,what are their respective ages?
Answer by mananth(16946)

Let Peter's age be x years
Arasoo's age = x+4
ali's age x-2 years
Their sum = 47 years
x+x+4+x-2=47
3x=45
x=15
Peter's age = 15
Arasoo's age = x+4 , =19 years
Ali's age = x-2 . = 13 years